# Does Object Binding Naturally Emerge in Large Pretrained Vision Transformers?

### v1

@@ 4.3 Attention weights (query-key similarity) correlate with IsSameObject
In Section 3.2 we showed that IsSameObject is best decoded quadratically. Since self-attention is also a quadratic interaction, binding information in the residual stream at layer $\ell$ can in principle guide where attention is allocated at layer $\ell+1$, allowing the model to selectively route attention within the same object to build a coherent object-level representation.
@@
  • Uninformed Ablation: Randomly shuffle $b(x_{i})$ across patches in the image at a specified ratio
  • Informed Ablation/Injection: Using ground-truth instance masks, we ablate or inject the true IsSameObject signal by linearly combining the mean object direction with each patch’s binding vector $b_{i}$: $\tilde{b}_{i}=(1-\alpha)\,\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}b_{\mathrm{object},j}+\alpha\,b_{\mathrm{object},i}.$
@@
Table 2: Ablations demonstrate the functional role of IsSameObject. Segmentation accuracy, instance accuracy, and DINO loss on layer 18 under uninformed (random shuffle) and informed (ground‑truth injection) ablations. Higher uninformed ratios mean more aggressive shuffling; lower informed $\alpha$ values mean stronger injection of true object labels.
@@
Segmentation Acc. (%) 44.14 41.03 39.20 44.14 44.91 43.59
Instance Acc. (%) 35.14 31.39 28.19 35.14 36.37 37.02
@@ 5 Limitations
We assume the trained probe cleanly splits each patch embedding into “feature” and “binding” components, a simplification that would benefit from further empirical exploration. Additionally, while randomizing the IsSameObject subspace clearly increases DINO’s pretraining loss, this indirect evidence falls short of demonstrating a causal link to downstream task performance or isolating its precise contribution. Finally, our downstream evaluations focus only on segmentation, leaving open whether these emergent binding signals also benefit other vision tasks such as visual reasoning. More broadly, this paper studies object-identity binding, i.e., whether two tokens belong to the same object; compositional binding, such as reusing object structure across contexts, is not explored here and is left for future work.
@@ 6 Conclusion
In this paper, we show that object binding naturally emerges in large, pretrained vision transformers, especially in DINOv2, and this effect is consistent across multiple models. We also show that it is an acquired rather than innate ability through comparisons with supervised models. IsSameObject, whether two patches belong to the same object, is reliably decodable and lies in a low-dimensional latent space. Ablating this signal both degrades downstream segmentation and opposes the model’s pretraining loss, suggesting that emergent object binding is a natural solution to the self‑supervised objective. Our study bridges what psychologists identify as object binding with emergent behavior in ViTs, challenges the belief that ViTs lack this ability and demonstrates how symbolic processing can arise in connectionist models.
Moving forward, we propose that AI researchers seeking to address binding failures adopt an alternative to explicit binding modules (such as Slot Attention [10]) by enhancing ViTs’ inherent object‑binding through tailored training objectives or minimal architectural tweaks. Future work could extend our approach to video models, in which object persistence across time is essential [11, 63]; such models should exhibit similar spatial binding seen in ViTs and additionally develop temporal binding to maintain object identity across frames. Also of interest will be how these object representations formed by binding interact with one another, with one possibility being that they interact with low-dimensional “object files” [64]. Together, these efforts will deepen our understanding of how symbolic processing of objects can emerge in connectionist models.
@@ A.1 Experimental Setup
All probes are trained for 16 epochs with a batch size of 256. For each sample image, we randomly select 64 patches and train with supervision.
For pairwise probes, we apply supervision over the upper triangular portion of the $64\times 64$ pairwise matrix. The model is optimized using binary cross-entropy loss.
@@
    $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cls}}+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}$ (3)
    where $\mathcal{L}_{\text{cls}}$ is replaced with a weighted binary cross-entropy loss distinguishing between object and no-object classes:
    $\mathcal{L}_{\text{cls}}=-w_{\text{obj}}\log(p_{\text{obj}})-w_{\text{no-obj}}\log(1-p_{\text{obj}})$ (4)
@@
    The hyperparameters follow the DETR configuration: no_object_weight = 0.1, mask_weight ($\lambda_{\text{mask}}$) = 5.0, dice_weight ($\lambda_{\text{dice}}$) = 5.0, and num_object_queries = 100.
A.2 Probe Performance
We train and evaluate quadratic probes on various ViTs including the complete DINOv2 family (small, base, large, giant) and CLIP-L14. All models utilize a patch size of 14×14 pixels. CLIP-L/14 processes 224×224 input images, resulting in 256 patches per image (16×16 grid).
Figure 7: Layer-wise IsSameObject classification accuracy across vision transformer architectures. DINOv2 models consistently achieve  88% accuracy, demonstrating robust development of IsSameObject representations across all model sizes. CLIP-L/14 reaches  85% accuracy and performs closer to the baseline level.
Overall, we observe that DINOv2 models achieve consistent high classification accuracy at 88%, and progress along layers, while CLIP model achieves a lower 84.8% and displays no such growing trend. DINO models seem to have object binding better than CLIP, although it is possible that less fine-grained grid leads to worse IsSameObject, since for each patch, the semantics are entangled.
A.3 Cross-Layer Binding
Cross-layer binding may actively happen under these conditions:
  • Layer-specific information processing: ViTs represent different types of information at different layers. Although the self-attention implements a "full range" interaction (all tokens to all tokens interaction), the model learns to limit the "interaction" between the correct tokens and integrate features of the object they belong to as we go deeper
  • Information retrieval from earlier layers: ViTs may access information from earlier layers when current layers lack necessary binding cues. As shown in Appendix A.4.3 positional information is gradually removed as the network deepens, and other low-level features may also be discarded. This information loss may trigger cross-layer retrieval as a computational strategy, which rather than maintaining all binding-relevant information in every layer, the model efficiently accesses what it needs from earlier representations. This hypothesis requires deeper analysis in future research.
Cross-layer interactions in ViTs can occur via the residual skip connections, so we also train quadratic probes between non-adjacent layers. Specifically, we compute $phi(x,y)=x^{T}W_{1}^{T}W2y$, where $W_{1}$ and $W_{2}$ are the learned projection matrices from layer 15 and layer 18, respectively. The probe accuracies are: layer 15: 89.0%, layer 18: 90.1%, layer 15-layer18: 83.3%, showing decent cross-layer object binding, and a certain degree of object binding across the system.
@@ A.4.1 Positional Encoding Distinguishes Identical Objects
A single transformer encoder layer can be viewed as comprising two complementary types of computations: token-wise (i.e., position-wise) operations, which act locally on individual tokens and can be executed in parallel to extract features and short-range interactions; and cross-token operations, implemented through scaled-dot-product attention, which enables long-range interactions by integrating contextual information across all tokens. When inspecting the mathematical formulation, we can also show that for identical tokens (i.e., identical patches), positional encoding is the only information that can guide the cross-token interactions. Here we review the operations for a transformer encoder with a single head from [65] in order. For simplicity, we also assume that query-key vectors have the same dimension as the model (i.e., $d_{k}=d$). We use blue color for the token-wise operations and red color for cross-token operations. For a sequence of input tokens $\mathbf{t}_{i}\in\mathbb{R}^{k}$ where $k=\text{n-channels}\times\text{patch-height}\times\text{patch width}$:
@@
$\quad\text{{\color[rgb]{0,0,1}embedding}}:\mathbf{e}_{i}=\mathbf{t}_{i}\mathbf{W}_{E}$
@@
$\quad\text{{\color[rgb]{0,0,1}Query-Key-Value}}:\mathbf{q}_{i}=\mathbf{x}_{i}\mathbf{W}_{Q},~\mathbf{k}_{i}=\mathbf{x}_{i}\mathbf{W}_{K},~\mathbf{v}_{i}=\mathbf{x}_{i}\mathbf{W}_{V}$
$\quad\text{{\color[rgb]{1,0,0}self-attention}}:\mathbf{U}=\text{softmax}\left(\frac{{\mathbf{Q}}(\mathbf{K})^{\intercal}}{\sqrt{d}}\right)\mathbf{V}$
$\quad\text{{\color[rgb]{0,0,1}projection MLP}}:\mathbf{y}_{i}=\mathbf{u}_{i}\mathbf{W}_{O}$
$\quad\text{{\color[rgb]{0,0,1}residual connection}}:\mathbf{y}_{i}=\mathbf{x}_{i}+\mathbf{y}_{i}$
$\quad\text{{\color[rgb]{0,0,1}normalization}}:\mathbf{z}_{i}=\text{LayerNorm}(\mathbf{y}_{i})$
$\quad\text{{\color[rgb]{0,0,1}feed-forward network}}:\mathbf{z}_{i}=\text{ReLU}(\mathbf{z}_{i}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}2$
$\quad\text{{\color[rgb]{0,0,1}residual connection}}:\mathbf{z}_{i}=\mathbf{z}_{i}+\mathbf{y}_{i}$
@@ A.4.2 Quantifying the Degree of Distinguishing Identical Objects
Figure 8: Layer-wise visualization of IsSameObject predictions on the test image (two identical red cars and one red boat.)

### v2

@@ 4.3 Attention weights (query-key similarity) correlate with IsSameObject
In Section 3.2 we showed that IsSameObject is best decoded quadratically. Since self-attention is also a quadratic interaction, binding information in the residual stream at layer $\ell$ can in principle guide how attention is allocated at layer $\ell+1$, allowing the model to selectively route attention within the same object to build a coherent object-level representation.
@@
  • Uninformed Ablation: Randomly shuffle $b(x_{i})$ across patches in the image at a specified ratio.
  • Informed Ablation (Injection): Using ground-truth instance masks, we inject the true IsSameObject signal by linearly combining the mean object direction with each patch’s binding vector $b_{i}$: $\tilde{b}_{i}=(1-\alpha)\,\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}b_{\mathrm{object},j}+\alpha\,b_{\mathrm{object},i}.$
@@
Table 2: Ablations demonstrate the functional role of IsSameObject. Segmentation mIoU, instance mIoU, and DINO loss on layer 18 under uninformed (random shuffle) and informed (ground‑truth injection) ablations.
@@
Segmentation mIoU (%) 44.14 41.03 39.20 44.14 44.91 43.59
Instance mIoU (%) 35.14 31.39 28.19 35.14 36.37 37.02
@@ 5 Limitations
We assume the trained probe cleanly splits each patch embedding into “feature” and “binding” components, a simplification that would benefit from further empirical exploration. We do not establish a causal relationship between object binding and downstream task performance, and further analysis is needed to understand how different pretraining objectives induce object binding. Finally, our downstream evaluations focus only on segmentation, leaving open whether these emergent binding signals also benefit other vision tasks such as visual reasoning. More broadly, this paper studies object binding at the patch level; more general forms of binding are not explored and are left for future work.
@@ 6 Conclusion
In this paper, we show that object binding naturally emerges in large, pretrained vision transformers, especially in DINOv2, and this effect is consistent across multiple models. We also show that it is an acquired rather than innate ability through comparisons across vision models. IsSameObject, whether two patches belong to the same object, is reliably decodable and lies in a low-dimensional latent space. Our results emergent object binding arises as a natural solution to self-supervised learning objectives. More broadly, our study bridges what psychologists identify as object binding with emergent behavior in ViTs, challenges the belief that ViTs lack such ability.
Looking ahead, we suggest that addressing binding failures in vision models may not require explicit object-centric modules (e.g., Slot Attention [29]), but could instead be achieved by strengthening the intrinsic object-binding mechanisms of ViTs through tailored training objectives or minimal architectural modifications. Another important direction for future work is to study how bound object representations interact with one another, potentially through low-dimensional “object files” [21]. Together, these efforts will deepen our understanding of how symbolic processing of objects can emerge in connectionist models.
@@ A.1 Experimental Setup
All probes are trained for 16 epochs with a batch size of 256 or 128.
A.2 Probe Details
Transformers propagate information across layers according to the above equations 12, where $h^{(\ell)}$ is the residual‑stream output. We use $h^{(\ell)}$ as the patch embedding for probing at layer $\ell$.
For pairwise probes, we apply supervision over the pairwise similarity matrix between all patches in an image. In practice, to reduce computational cost, we randomly sample 64 patches per image in each epoch and apply supervision only to the resulting patch pairs. The probes are trained using binary cross-entropy loss.
@@ A.2 Probe Details
    $\mathcal{L}_{\text{total}}=\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}$ (3)
@@
    The hyperparameters follow the DETR configuration: mask_weight ($\lambda_{\text{mask}}$) = 5.0, dice_weight ($\lambda_{\text{dice}}$) = 5.0. Since only 64 patches are sampled per image in each epoch, we use a reduced number of object queries, num_object_queries = 10.
A.2.1 Quadratic probe details
We enforce symmetry in $W$ to reflect the property $\textit{IsSameObject}(x,y)=\textit{IsSameObject}(y,x)$. All quadratic probes apply a $1/\sqrt{C}$ normalization. In the full quadratic probe, $\ell_{2}$ weight decay encourages a low effective rank. The fixed-rank probe instead explicitly upper-bounds $\mathrm{rank}(W)$ through its factorized construction. For analysis, we recover $W1$ and $W2$ from the expression $\sigma\!\left(x^{\top}W_{1}^{\top}W_{2}\,y+b\right)$ in Equation 3.1, by computing the singular value decomposition (SVD) of the learned symmetric matrix $W$.
We do not directly parameterize the model using a fixed-rank $W_{1}$ together with a sign matrix $S$ that defines $W_{2}$ by flipping rows of $W_{1}$, as such a discrete sign matrix is difficult to optimize in a fully differentiable manner.
Algorithm 1 Quadratic Probe (full rank)
[…]
1: Learn $A\in\mathbb{R}^{C\times C}$, $b\in\mathbb{R}$
[…]
4:return $\sigma(s)$
Algorithm 2 Quadratic Probe (with fixed rank $r$)
[…]
1: Learn $U,V\in\mathbb{R}^{r\times C}$, $b\in\mathbb{R}$
[…]
4:return $\sigma(s)$
A.3 Probe Performance
A.3.1 Baselines
We consider several baselines for the IsSameObject task.
Majority baseline.
As a trivial baseline, we always predict that a patch pair belongs to different objects, reflecting the strong class imbalance in the ADE20K dataset. This baseline achieves an accuracy of 72.6%.
Distance-based baseline.
We also include a distance-based baseline that captures dataset statistics. A single scalar threshold is learned on the pairwise patch embedding distance: if the distance is below the threshold, the pair is predicted to belong to the same object; otherwise, it is predicted to be different. This baseline achieves an accuracy of 77.16%.
Similarity- and attention-based probes.
We further compare our IsSameObject probes against methods motivated by prior work on feature similarity and attention-based object selectivity. For all baselines in this category, only a scalar bias term is trained, while the underlying representations are kept fixed. Specifically, we consider:
  • Cosine similarity probe: $\textit{IsSameObject}(x,y)=\sigma\!\left(-\frac{x^{\top}y}{\|x\|_{2}\,\|y\|_{2}}+b\right)$;
  • Dot-product probe: $\textit{IsSameObject}(x,y)=\sigma(-x^{\top}y+b)$;
  • Self-attention probe: $\textit{IsSameObject}(x,y)=\sigma(-x^{\top}W_{\text{query}}W_{\text{key}}\,y+b)$, where $W_{\text{query}}$ and $W_{\text{key}}$ are the query and key projection matrices from the self-attention layer at layer $\ell+1$.
These results (Fig.8) demonstrate that quadratic probes capture IsSameObject structure beyond what can be explained by feature similarity or attention alone.
Figure 8: Layer-wise IsSameObject baseline accuracy in DINOv2-Large.
A.3.2 Cross-Model Comparison
We train and evaluate quadratic probes on a range of Vision Transformers (ViTs), including the full DINOv2 family (Small, Base, Large, and Giant), as well as CLIP, MAE, and an ImageNet-supervised ViT-Large model. The evaluated backbones are listed in Table 3.
Table 3: Vision Transformer backbones used for evaluation.
Model family HuggingFace identifier
DINOv2-Small facebook/dinov2-small
DINOv2-Base facebook/dinov2-base
DINOv2-Large facebook/dinov2-large
DINOv2-Giant facebook/dinov2-giant
CLIP ViT-L/14 openai/clip-vit-large-patch14
MAE ViT-L facebook/vit-mae-large
ViT-L (IN1K) google/vit-large-patch16-224
All DINOv2 models use a patch size of $14\times 14$ pixels and operate on raw inputs of size $518\times 518$. To enable fair comparison across models with different native input resolutions and patch sizes, we standardize evaluation by matching DINOv2’s per-patch spatial coverage. Specifically, inputs to all other models are resized such that their patch grids align with those of DINOv2, resulting in identical spatial patch divisions and, consequently, the same majority baseline accuracy (72.6%).
For example, CLIP ViT-L/14 natively processes $224\times 224$ images. To match DINOv2’s patch grid, the input image is first resized so that its shortest edge is 518 pixels (as in DINOv2), then cropped into multiple $224\times 224$ regions, each of which is patchified using $14\times 14$ patches. This procedure ensures consistent spatial correspondence across models despite differences in architecture and pretraining.
Figure 9: IsSameObject predictions distinguish objects in complex scenes. Additional visualizations of quadratic probe results for DINOv2-Large at layer 18. The deer image is taken from the DINOv3 paper [49].
Figure 10: Layer-wise IsSameObject classification accuracy across ViT-Large under different pretraining objectives. DINOv2 achieves the highest performance, exceeding 90% accuracy in upper layers, while MAE remains close to the baseline, indicating little to no emergent object binding.
Figure 11: Layer-wise IsSameObject classification accuracy across the DINOv2 family as a function of model size. DINOv2-Large achieves the strongest overall performance, while peak accuracy is comparable across model scales, reaching approximately 88%.
In Fig. 10, we observe that for ViT-Large models with different pretraining objectives, object-binding performance in DINOv2 begins to emerge in the middle layers and peaks in the later layers, whereas ImageNet-supervised ViT and CLIP remain largely consistent across layers.
In Fig. 11, we observe that within the DINOv2 family, smaller models tend to develop strong object-binding representations at later normalized layers.
@@ A.4.1 Positional Encoding Distinguishes Identical Objects
A single transformer encoder layer can be viewed as comprising two complementary types of computations: token-wise (i.e., position-wise) operations, which act locally on individual tokens and can be executed in parallel to extract features and short-range interactions; and cross-token operations, implemented through scaled-dot-product attention, which enables long-range interactions by integrating contextual information across all tokens.
When inspecting the mathematical formulation, we can also show that for identical tokens (i.e., identical patches), positional encoding is the only information that can guide the cross-token interactions. Here we review the operations for a transformer encoder with a single head from [58] in order. For simplicity, we also assume that query-key vectors have the same dimension as the model (i.e., $d_{k}=d$). We use blue color for the token-wise operations and red color for cross-token operations. For a sequence of input tokens $\mathbf{t}_{i}\in\mathbb{R}^{k}$ where $k=\text{n-channels}\times\text{patch-height}\times\text{patch width}$:
@@
$\quad\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}embedding}}:\mathbf{e}_{i}=\mathbf{t}_{i}\mathbf{W}_{E}$
@@
$\quad\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Query-Key-Value}}:\mathbf{q}_{i}=\mathbf{x}_{i}\mathbf{W}_{Q},~\mathbf{k}_{i}=\mathbf{x}_{i}\mathbf{W}_{K},~\mathbf{v}_{i}=\mathbf{x}_{i}\mathbf{W}_{V}$
$\quad\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}self-attention}}:\mathbf{U}=\text{softmax}\left(\frac{{\mathbf{Q}}(\mathbf{K})^{\intercal}}{\sqrt{d}}\right)\mathbf{V}$
$\quad\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}projection MLP}}:\mathbf{y}_{i}=\mathbf{u}_{i}\mathbf{W}_{O}$
$\quad\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}residual connection}}:\mathbf{y}_{i}=\mathbf{x}_{i}+\mathbf{y}_{i}$
$\quad\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}normalization}}:\mathbf{z}_{i}=\text{LayerNorm}(\mathbf{y}_{i})$
$\quad\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}feed-forward network}}:\mathbf{z}_{i}=\text{ReLU}(\mathbf{z}_{i}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}2$
$\quad\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}residual connection}}:\mathbf{z}_{i}=\mathbf{z}_{i}+\mathbf{y}_{i}$
@@
Proposition A.1: Positional Encoding Breaks Symmetry For a transformer encoder layer, if two input tokens satisfy $\mathbf{t}_{i}=\mathbf{t}_{j}$ and $\mathbf{p}_{i}=\mathbf{p}_{j}$, then their self-attention outputs are identical, i.e., $\mathbf{u}_{i}=\mathbf{u}_{j}$. Therefore, for identical tokens, positional encoding is the only signal that can differentiate their cross-token interactions.
@@ A.4.2 Quantifying the Degree of Distinguishing Identical Objects
Figure 12: Layer-wise visualization of IsSameObject predictions on the test image with two identical red cars and one red boat.